A block of mass m =12kg is pushed at a constant speed up an incline plane angle of inclination 30 though a distance of s =4.5mand raised to a height h=2m the coffiepnt of kinetic fiction between the block and the incline plane is 0.41 calculate the work done by normal force the external force applied parallel to the incline gravity the frictional force

1
Expert's answer
2021-12-14T15:09:14-0500

The work done by normal force:

"W=mg\cos\theta·h=36\text{ J}."

The work done by the frictional force:


"W=\mu mg\cos\theta·h=15\text{ J}."

The work done by the external force:


"W=\mu mg\cos\theta·h=15\text{ J}."